OpenStudy (anonymous):

find the specified term of each geometric sequence. a1=9,r=1/3;a5

OpenStudy (tkhunny):

Didn't we just do this? a1 = 9 a2 = 9*(1/3) a3 = 9*(1/3)^2 etc.

OpenStudy (bradely):

a5 =a1r^5-1 =9*(1/3)^4 =1/9 Source: http://www.mathskey.com/question2answer/